OpenStudy (anonymous):

Factor x2 - x - 90 = 0

OpenStudy (anonymous):

POOOOOF BISHHH

OpenStudy (unklerhaukus):

\[x^2-x-90=(x+a)(x+b)=x^2+(a+b)x+ab\] \[a+b=-1,\qquad ab=-90\]